# Low-rank orthogonalization for large-scale matrix optimization with applications to foundation model training

### v1

@@ Abstract
Neural network (NN) training is inherently a large-scale matrix optimization problem, yet the matrix structure of NN parameters has long been overlooked. Recently, the optimizer Muon [28], which explicitly exploits this structure, has gained significant attention for its strong performance in foundation model training. A key component contributing to Muon’s success is matrix orthogonalization. In this paper, we propose low-rank orthogonalization, which explicitly leverages the low-rank nature of gradients during NN training. Building on this, we propose low-rank matrix-signed gradient descent and a low-rank variant of Muon. Our numerical experiments demonstrate the superior performance of low-rank orthogonalization, with the low-rank Muon achieving promising results in GPT-2 and LLaMA pretraining—surpassing the performance of the carefully tuned vanilla Muon. Theoretically, we establish the iteration complexity of the low-rank matrix-signed gradient descent for finding an approximate stationary solution, as well as that of low-rank Muon for finding an approximate stochastic stationary solution under heavy-tailed noise.
@@ 1 Introduction
Training neural networks (NNs) [32], particularly recent foundation models [8], has consistently posed challenging large-scale optimization problems. Over the past decade, NN training has been dominated by vector-variate optimization methods—including SGD [9], AdaGrad [18], RMSprop [25], Adadelta [55], Adam [29], and AdamW [37]. Nonetheless, these methods disregard the inherent matrix structure of NN parameters—such as those in multi-layer perceptrons [45], convolutional layers [33], and the query, key, and value projections in attention mechanisms [52].
Recently, a shift has taken place: optimization methods that exploit matrix structure are receiving increasing attention and have begun to demonstrate strong performance in foundation model training [28, 35, 41]. These methods focus on solving the matrix optimization problem:
@@
In particular, Shampoo, developed in [5, 21], applies left and right preconditioning matrices as follows:
@@
where $\eta_{k}>0$ is the step size, $G^{k}\in\mathbb{R}^{m\times n}$ denotes the (stochastic) gradient of $f$ at $X^{k}$, and $L^{k}\in\mathbb{R}^{m\times m}$ and $R^{k}\in\mathbb{R}^{n\times n}$ are the left and right preconditioning matrices, respectively. Shampoo updates the left and right preconditioners $\{L^{k}\}$ and $\{R^{k}\}$ using the second-order statistics of the accumulated gradients, similarly to AdaGrad, and has demonstrated comparable performance to popular vector-variate optimizers such as Adam and AdamW on foundation model training. Following Shampoo, other matrix-variate optimizers that use two-sided preconditioners, such as CASPR [19] and SOAP [53], were also developed. A preconditioned Riemannian gradient descent method was developed in [7] for low-rank matrix recovery, which adopts only the diagonal part of the Shampoo preconditioners. Moreover, one-sided preconditioned variants of Shampoo were developed in [4, 54].
In addition to Shampoo and its variants, another matrix-variate optimizer, Muon [28], has attracted significant attention for outperforming standard optimizers such as Adam and AdamW in foundation model training [3, 35]. At each iteration, Muon performs the update:
@@
where $G(\cdot;\xi)$ is the stochastic gradient of $f(\cdot)$, and $\mathrm{msgn}(M^{k})=U^{k}(V^{k})^{T}$ denotes the matrix sign of $M^{k}$, with $U^{k}$ and $V^{k}$ being the left and right singular vectors of $M^{k}$, respectively. The matrix sign computation is often referred to as matrix orthogonalization, because calculating $\mathrm{msgn}(M)$ is equivalent to finding the (semi-)orthogonal matrix closest to $M$ with respect to the Frobenius norm (see, e.g., [6, Proposition 4]). Muon’s empirical success has sparked significant research interest, including efforts to understand its relationship with other algorithms, establish its convergence guarantees, and propose new variants (see, e.g., [4, 14, 15, 20, 30, 31, 34, 36, 38, 41, 43, 46, 47, 48]). A popular interpretation of Muon is from the perspective of a linear minimization oracle with respect to the spectral norm (e.g., see [6, 14, 20, 30, 31, 43]). That is, the matrix sign computation in (3) can be recast as:
$-\mathrm{msgn}(M^{k})=\underset{\|\Delta\|\leq 1}{\arg\min}\{\langle M^{k},\Delta\rangle\},$
where $\|\cdot\|$ denotes spectral norm. Based on this interpretation, algorithmic designs leveraging general matrix-induced norms $\|\cdot\|_{p\to q}$ have been discussed in [6, 14, 20, 43]. In addition, Muon is also connected to earlier algorithms and can be viewed as a special case of Shampoo, despite not explicitly using preconditioners. As discussed in [28], by taking $L^{k}=G^{k}(G^{k})^{T}$ and $R^{k}=(G^{k})^{T}G^{k}$ in (2), the Shampoo updates reduce to the matrix-signed update: $X^{k+1}=X^{k}-\eta_{k}\mathrm{msgn}(G^{k})$. Furthermore, convergence guarantees for Muon have been extensively studied (e.g., see [4, 15, 30, 34, 43, 46, 47, 48]), and numerous new variants—such as SWAN [38], Scion [41], Gluon [43], PolarGrad [31], Dion [1, 2], and AdaMuon [49]—have been proposed.
@@
Beyond applying orthogonalization, a key innovation of Muon is the use of a more GPU-friendly method—specifically, Newton-Schulz iterations (typically with five steps)—to perform inexact orthogonalization, which makes Muon well-suited for modern foundation model training. In fact, several earlier methods, including spectral gradient descent [10, 11, 12] and orthogonalized gradient descent [51], have applied orthogonalization using SVD to matrix optimization problems. However, since SVD is not computationally efficient in GPU environments, these methods fail to scale to foundation model training.
@@
Inspired by Muon and its orthogonalization subroutine, we aim at developing a faster, lightweight orthogonalization method to be implemented in Muon and its variants. Specifically, our design leverages the widely observed phenomenon that the gradient matrices of NN parameters are often low-rank (see, e.g., [23, 26, 39, 57]). To exploit this low-rank property, we propose performing low-rank orthogonalization by incorporating well-known matrix approximation techniques for low-rank matrices [16, 22]. Our approach first constructs a low-rank projection of the gradient matrix using QR decomposition on a sketched matrix, and then performs orthogonalization on the projected matrix by leveraging its structure. Our proposed low-rank orthogonalization offers two main advantages over common orthogonalization:
@@
  • Computational efficiency: Orthogonalization can be seen as computing the polar factor of a given full matrix. In contrast, our low-rank orthogonalization first computes the Q factor of a smaller sketched matrix, followed by the polar factor of a projected matrix constructed using the Q factor. Since both the QR decomposition and polar decomposition are performed on much smaller matrices, our low-rank approach enjoys substantial computational savings for large-scale problems.
@@
  • Noise robustness: In the presence of noise, singular vectors associated with small singular values often vary significantly, leading to instability in orthogonalization via Newton–Schulz iterations. By contrast, our low-rank orthogonalization method clips these singular vectors to eliminate unstable estimates of singular vectors associated with small singular values, thereby stabilizing the orthogonalization process and yielding a robust estimate of the matrix sign.
@@
These advantages will be illustrated in detail in Sections 3.1 and 4. Based on low-rank orthogonalization, we develop low-rank matrix-signed gradient descent and a low-rank variant of Muon. We also establish their complexity guarantees under mild assumptions.
@@
  • We propose low-rank orthogonalization to be incorporated into matrix-variate optimization algorithms such as Muon. It can be efficiently executed on GPUs and serves as a lightweight substitute for existing orthogonalization methods such as Newton-Schulz iterations.
  • Under mild assumptions, we establish the iteration complexity of low-rank matrix-signed gradient descent and a low-rank variant of Muon, respectively. To the best of our knowledge, our complexity analysis of low-rank Muon provides the first result for Muon-type algorithms—including vanilla Muon—under heavy-tailed noise.
The remainder of this paper is organized as follows. In Section 2, we introduce the notation and assumptions used throughout the paper. In Section 3, we propose low-rank orthogonalization and, based on it, develop low-rank matrix-sign gradient descent and low-rank Muon. In Section 4, we present numerical results. Finally, we provide the proofs of the main results and concluding remarks in Sections 5 and 6, respectively.
@@ 2 Notation and assumptions
Assumption 1(a) is standard. Assumption 1(b) is natural in the analysis of Muon-type algorithms and has been used previously (e.g., see [15, 43, 48]). It follows from Assumption 1(b) that
@@ 3 Matrix optimization with low-rank orthogonalization
In this section, we propose algorithms with low-rank orthogonalization for solving (1). In particular, we first propose low-rank orthogonalization in Section 3.1, which serves as a subroutine in matrix-variate optimization algorithms. Then, we propose low-rank matrix-signed gradient descent methods in Section 3.2, and a low-rank variant of Muon in Section 3.3.
@@ 3.1 Low-rank orthogonalization
Orthogonalization has attracted increasing attention in recent optimizer designs, as it has shown strong empirical promise in foundation model training (e.g., see [6, 28, 31, 51]). In this subsection, we develop a low-rank orthogonalization method, leveraging low-rank matrix approximation techniques, that serves as a lightweight substitute for the orthogonalization subroutine used in matrix-variate optimizers.
Specifically, our low-rank orthogonalization method, presented in Algorithm 1, is based on Gaussian sketching [22]. This method first draws a Gaussian random matrix $G\in\mathbb{R}^{n\times r}$ with $r\ll\varrho$, and performs a QR decomposition on $MG$ to obtain a column-orthogonal Q factor $Q\in\mathbb{R}^{m\times r}$. Then, it computes $\mathrm{msgn}(Q^{T}M)\in\mathbb{R}^{r\times n}$ and returns $M_{O}=Q\mathrm{msgn}(Q^{T}M)$ as a low-rank approximation for $\mathrm{msgn}(M)$. As will be shown in Theorem 1, $M_{O}$ represents the matrix sign of $QQ^{T}M$, which is a low-rank approximation of $M$. Its proof is deferred to Section 5.1.
@@
Moreover, we have
$\mathrm{msgn}(QQ^{T}M)=Q\mathrm{msgn}(Q^{T}M).$ (6)
@@
The relation (5) is adapted from [22, Theorem 10.5], where additional guarantees—such as those involving different matrix norms and those providing high-probability bounds—can also be found. In addition, low-rank matrix approximation based on column selection (e.g., see [16, 17]) can also be used to develop a low-rank orthogonalization method. However, since its approximation guarantee is more complicated than that of the Gaussian sketching-based approach, we defer the column-selection-based low-rank orthogonalization method to Algorithm 5 in Appendix A.
@@
Computational efficiency:
We compare the computation time on GPUs for calculating inexact matrix sign of high-dimensional matrices using Newton-Schulz iterations, low-rank orthogonalization based on Gaussian sketching (Algorithm 1) and column selection (Algorithm 5), and truncated SVD.
For each $n\in\{1000,2000,5000,10000\}$, we generate 50 random matrices $M\in\mathbb{R}^{n\times n}$, with each entry following the standard Gaussian distribution. We then apply all competing orthogonalization methods to estimate $\mathrm{msgn}(M)$. We implement Newton-Schulz iterations as provided in Muon [28] with 5 iterations. For our low-rank orthogonalization methods, we set $r=0.1n$ as the input rank parameter, we use command torch.linalg.qr to compute the Q factor, and apply 5 iterations of Newton-Schulz scheme following Muon [28] to compute the matrix sign of $Q^{T}M$. In addition, we use command torch.pca_lowrank for efficiently performing truncated SVD, setting the rank to $0.1n$ for the truncation.
@@
We present a comparison of the computation time on GPUs for all competing methods in Figure 1, and the time distribution of our low-rank orthogonalization method with Gaussian sketching in Figure 2. From Figure 1, we observe that our low-rank orthogonalization method significantly reduces computation time compared to Newton-Schulz iterations and truncated SVD. Although both truncated SVD and our low-rank orthogonalization exploit low-rank structure for computation, truncated SVD is not well-suited to GPU environments and is therefore slower than the Newton-Schulz iterations. From Figure 2, we observe that in our low-rank orthogonalization method with Gaussian sketching, the QR decomposition accounts for approximately half of the total time, while the Newton-Schulz iterations and other operations (such as generating Gaussian random matrices and performing matrix multiplications) each make up roughly half of the remaining time.
Noise robustness:
@@
For each $n\in\{1000,2000,5000,10000\}$, we first randomly generate 10 nearly low-rank matrices $M\in\mathbb{R}^{n\times n}$ following strategy: the top $0.1n$ singular values are set to $1$, the remaining singular values are set to $10^{-4}$, and the singular vectors are randomly generated orthogonal vectors. For each $M$, we generate 50 noise matrices $N\in\mathbb{R}^{n\times n}$, with each entry drawn from a Gaussian distribution with mean zero and variance $\sigma^{2}\in\{0.1,1,10\}$, and construct noisy matrices $M_{N}=M+N$. We next apply Newton-Schulz iterations, and our low-rank orthogonalization method based on Gaussian sketching (Algorithm 1), using an input rank parameter $0.1n$, to estimate $\mathrm{msgn}(M_{N})$. [^1] We omit low-rank orthogonalization with column selection (Algorithm 5) because its performance is similar to that of the version with Gaussian sketching. For each $M$, we compute the trace of the empirical covariance matrix of the estimates of $\mathrm{msgn}(M_{N})$ produced by both Newton-Schulz iterations and Algorithm 1. Both methods are implemented in the same way as in the above synthetic experiments to illustrate the computational efficiency.
@@
Figure 3 presents a comparison of the average trace of the covariance matrices across all competing methods for matrix sign estimation. From this figure, we observe that compared to applying the Newton-Schulz iterations to the full matrix, our low-rank orthogonalization method yields matrix sign estimates that are significantly less sensitive to noise. This is because singular vectors associated with small singular values are more sensitive to noise than those corresponding to large singular values. As a result, our low-rank method provides a more stable and robust estimate of the matrix sign function, particularly for nearly low-rank matrices.

### v2

# Low-Rank Orthogonalization for Large-Scale Matrix Optimi- zation with Applications to Foundation Model Training

@@ Abstract
Neural network (NN) training is inherently a large-scale matrix optimization problem, yet the matrix structure of NN parameters has long been overlooked. Recently, the optimizer Muon (Jordan et al., ), which explicitly exploits this structure, has gained significant attention for its strong performance in foundation model training. A key component contributing to Muon’s success is matrix orthogonalization. In this paper, we propose low-rank orthogonalization, which performs orthogonalization by leveraging the low-rank nature of gradients during NN training. Building on this, we introduce low-rank matrix-signed gradient descent (MSGD) and a low-rank variant of Muon. Numerical experiments demonstrate the superior performance of low-rank orthogonalization, with low-rank Muon achieving promising results in GPT-2 and LLaMA pretraining—surpassing the carefully tuned vanilla Muon on tasks with large model sizes. Theoretically, we establish the iteration complexity of low-rank MSGD for finding an approximate stationary solution, and the iteration complexity of low-rank Muon for finding an approximate stochastic stationary solution under heavy-tailed noise. The code to reproduce our numerical experiments is available at https://github.com/dengzhanwang/Low-rank-Muon.
@@ 1 Introduction
Training neural networks (NNs) (LeCun et al., 2015), particularly recent foundation models, has consistently posed challenging large-scale optimization problems. Over the past decade, NN training has been dominated by vector-variate optimization methods—including SGD (Bottou, 2010), AdaGrad (Duchi et al., 2011), RMSprop (Hinton et al., 2012), Adadelta (Zeiler, 2012), Adam (Kingma and Ba, 2015), and AdamW (Loshchilov and Hutter, 2019). Nonetheless, these methods disregard the inherent matrix structure of NN parameters—such as those in multi-layer perceptrons (Rosenblatt, 1958), convolutional layers (LeCun et al., 1989), and the query, key, and value projections in attention mechanisms (Vaswani et al., 2017).
Recently, a shift has occurred: optimization methods that exploit matrix structure are receiving increasing attention and have begun to demonstrate strong performance in foundation model training (Jordan et al., ; Pethick et al., 2025). These methods focus on solving the matrix optimization problem:
@@
In particular, Shampoo, as developed in (Anil et al., 2020; Gupta et al., 2018), applies left and right preconditioning matrices and updates them as follows:
@@
where $\eta_{k}>0$ is the step size, $G^{k}\in\mathbb{R}^{m\times n}$ denotes the (stochastic) gradient of $f$ at $X^{k}$, and $L^{k}\in\mathbb{R}^{m\times m}$ and $R^{k}\in\mathbb{R}^{n\times n}$ are the left and right preconditioning matrices, respectively. Shampoo updates the left and right preconditioners $\{L^{k}\}$ and $\{R^{k}\}$ using the second-order statistics of the accumulated gradients, similarly to AdaGrad, and has shown comparable performance to popular vector-variate optimizers such as Adam and AdamW on foundation model training. Following Shampoo, other matrix-variate optimizers that use two-sided preconditioners, such as CASPR (Duvvuri et al., 2024) and SOAP (Vyas et al., 2025), were also developed. A preconditioned Riemannian gradient descent method was developed in (Bian et al., 2024) for low-rank matrix recovery, which adopts only the diagonal part of the Shampoo preconditioners. Moreover, one-sided preconditioned variants of Shampoo were developed in (An et al., 2025; Xie et al., 2025).
In addition to Shampoo and its variants, another matrix-variate optimizer, Muon (Jordan et al., ), has attracted significant attention for outperforming standard optimizers such as Adam and AdamW in foundation model training (AI et al., 2025). At each iteration, Muon performs the update:
@@
where $G(\cdot;\xi)$ denotes the stochastic gradient of $f(\cdot)$, and $\mathrm{msgn}(M^{k})=U^{k}(V^{k})^{T}$ denotes the matrix sign of $M^{k}$, with $U^{k}$ and $V^{k}$ containing the left and right singular vectors of $M^{k}$, respectively. The matrix sign computation is often referred to as matrix orthogonalization, because calculating $\mathrm{msgn}(M)$ is equivalent to finding the (semi-)orthogonal matrix closest to $M$ with respect to the Frobenius norm (see, e.g., (Bernstein and Newhouse, 2024, Proposition 4)). Muon’s empirical success has sparked significant research interest, including efforts to understand its relationship with other algorithms, establish its convergence guarantees, and propose new variants (see, e.g., (An et al., 2025; Cesista, 2025; Chen et al., 2025; Glentis et al., 2025; Kovalev, 2025; Lau et al., 2025; Li and Hong, 2025; Liu et al., 2025; Ma et al., 2024; Pethick et al., 2025; Riabinin et al., 2025; Sato et al., 2025; Sfyraki and Wang, 2025; Shen et al., 2025)). A popular interpretation of Muon is from the perspective of a linear minimization oracle with respect to the spectral norm (e.g., see (Bernstein and Newhouse, 2024; Cesista, 2025; Glentis et al., 2025; Kovalev, 2025; Lau et al., 2025; Riabinin et al., 2025)). That is, the matrix sign computation in (3) can be recast as $\mathrm{msgn}(M^{k})={\arg\min}_{\|\Delta\|\leq 1}\{\langle M^{k},\Delta\rangle\}$, where $\|\cdot\|$ denotes spectral norm. Based on this interpretation, algorithmic designs leveraging general matrix-induced norms $\|\cdot\|_{p\to q}$ have been discussed in (Bernstein and Newhouse, 2024; Cesista, 2025; Glentis et al., 2025; Riabinin et al., 2025). In addition, Muon is also connected to earlier algorithms and can be viewed as a special case of Shampoo, despite not explicitly using preconditioners. As discussed in (Jordan et al., ), by taking $L^{k}=G^{k}(G^{k})^{T}$ and $R^{k}=(G^{k})^{T}G^{k}$ in (2), the Shampoo updates reduce to the matrix-signed update: $X^{k+1}=X^{k}-\eta_{k}\mathrm{msgn}(G^{k})$. Furthermore, convergence guarantees for Muon have been extensively studied (e.g., see (An et al., 2025; Chen et al., 2025; Kovalev, 2025; Li and Hong, 2025; Riabinin et al., 2025; Sato et al., 2025; Sfyraki and Wang, 2025; Shen et al., 2025)), and numerous new variants—such as SWAN (Ma et al., 2024), Scion (Pethick et al., 2025), Gluon (Riabinin et al., 2025), PolarGrad (Lau et al., 2025), Dion (Ahn and Xu, 2025; Ahn et al., 2025), and AdaMuon (Si et al., 2025)—have been proposed.
@@
Beyond applying orthogonalization, a key innovation of Muon is its use of a GPU-friendly method—Newton-Schulz iterations (typically with five steps)—to perform inexact matrix orthogonalization, making it well-suited for modern foundation model training. In fact, several earlier methods, including spectral gradient descent (Carlson et al., 2015a; c; b) and orthogonalized gradient descent (Tuddenham et al., 2022), have already adopted SVD-based orthogonalization for matrix optimization problems. However, since SVD is not computationally efficient on GPUs when training large-scale neural networks, these methods fail to scale to foundation model training.
@@
Inspired by Muon and its orthogonalization subroutine, we aim to develop a faster and more lightweight orthogonalization method to further enhance Muon and its variants. Specifically, our design leverages the widely observed phenomenon that the gradient matrices of NN parameters are often low-rank (see, e.g., (Hao et al., 2024; Hu et al., 2022; Malladi et al., 2023; Zhao et al., 2024)). To exploit this low-rank property, we propose performing low-rank orthogonalization by incorporating well-known low-rank matrix approximation techniques (Drineas et al., 2006a; Halko et al., 2011). Our approach first constructs a low-rank projection of the gradient matrix using QR decomposition on a sketched matrix, and then performs orthogonalization on the projected matrix by leveraging its structure. Our proposed low-rank orthogonalization offers two main advantages over existing orthogonalization methods:
  1. 1.
@@
    Computational Efficiency: Traditional orthogonalization can be seen as computing the polar factor of a given full matrix. To exploit the low-rank property, our low-rank orthogonalization first computes the $Q$ factor of a smaller sketched matrix, and then computes the polar factor of a projected matrix constructed using the $Q$ factor. Since both the QR decomposition and polar decomposition are performed on much smaller matrices, our low-rank approach enjoys substantial computational savings for large-scale problems.
  2. 2.
@@
    Noise Robustness: In the presence of noise, singular vectors associated with small singular values often vary significantly, leading to instability when directly applying orthogonalization to the full matrix. To circumvent this instability, our low-rank orthogonalization method clips these vectors to eliminate unreliable estimates, thereby stabilizing the orthogonalization process and yielding a robust estimate of the matrix sign.
@@
These advantages will be illustrated in detail in Sections 3.1 and 4. Based on low-rank orthogonalization, we develop low-rank matrix-signed gradient descent (MSGD) and a low-rank variant of Muon. We also establish their complexity guarantees under mild assumptions.
@@
  • We propose low-rank orthogonalization that can be readily incorporated into matrix-variate optimization algorithms such as Muon. It can be efficiently executed on GPUs and serves as a lightweight substitute for traditional orthogonalization methods that are directly applied to the full matrix, such as Newton-Schulz iterations.
  • Under mild assumptions, we establish the iteration complexity of low-rank MSGD and a low-rank variant of Muon. To the best of our knowledge, our results provide the first complexity guarantees for a broad class of inexact Muon-type algorithms, including vanilla Muon, under heavy-tailed noise.
The remainder of this paper is organized as follows. In Section 2, we introduce the notation and assumptions used throughout the paper. In Section 3, we propose low-rank orthogonalization and, based on it, develop low-rank MSGD and low-rank Muon. In Section 4, we present numerical results. Finally, we provide the proofs of the main results in Section 5.
@@ 2 Notation and Assumptions
Assumption 1(a) is standard. Assumption 1(b) is natural in the analysis of Muon-type algorithms (e.g., see (Chen et al., 2025; Riabinin et al., 2025; Shen et al., 2025)). It follows from Assumption 1(b) that
@@ 3 Matrix Optimization with Low-Rank Orthogonalization
In this section, we propose matrix optimization algorithms with low-rank orthogonalization for solving (1). In particular, we first propose low-rank orthogonalization in Section 3.1. Then, we propose low-rank MSGD in Section 3.2, and a low-rank variant of Muon in Section 3.3.
@@ 3.1 Low-Rank Orthogonalization
Orthogonalization techniques have attracted increasing attention in recent optimizer designs, as it has shown strong empirical performance in foundation model training (e.g., see (Bernstein and Newhouse, 2024; Jordan et al., ; Lau et al., 2025; Tuddenham et al., 2022)). In this subsection, we develop a low-rank orthogonalization method, leveraging low-rank matrix approximation techniques, that serves as a lightweight substitute for the existing orthogonalization subroutines used in matrix-variate optimizers.
@@
Specifically, our low-rank orthogonalization method, presented in Algorithm 1, is based on Gaussian sketching (Halko et al., 2011). This method first draws a Gaussian random matrix $G\in\mathbb{R}^{n\times r}$ with $r\ll\varrho$, and performs a QR decomposition on $MG$ to obtain a column-orthogonal Q factor $Q\in\mathbb{R}^{m\times r}$. Then, it computes $\mathrm{msgn}(Q^{T}M)\in\mathbb{R}^{r\times n}$ and returns $M_{O}=Q\mathrm{msgn}(Q^{T}M)$ as a low-rank approximation for $\mathrm{msgn}(M)$. As will be shown in Theorem 1, $M_{O}$ represents the matrix sign of $QQ^{T}M$, which is a low-rank approximation of $M$. Its proof is deferred to Section 5.1.
@@
Moreover, we have $\mathrm{msgn}(QQ^{T}M)=Q\mathrm{msgn}(Q^{T}M)$.
@@
The relation (5) is adapted from (Halko et al., 2011, Theorem 10.5), where additional guarantees—such as those involving different matrix norms and high-probability bounds—can also be found. In addition, low-rank matrix approximation based on column selection (e.g., see (Drineas et al., 2006a; b)) can be used to develop a low-rank orthogonalization method. However, since its approximation guarantee is more complicated than that of the Gaussian sketching-based approach, we defer the column-selection-based method to the supplementary materials.
@@
Computational Efficiency
We compare the computation time on GPUs for calculating inexact matrix sign of high-dimensional matrices using Newton-Schulz iterations, low-rank orthogonalization based on Gaussian sketching (Algorithm 1) and column selection (see supplementary materials), and truncated SVD.
For each $n\in\{1000,2000,5000,10000\}$, we generate 50 random matrices $M\in\mathbb{R}^{n\times n}$, with each entry drawn from the standard Gaussian distribution. We then apply all competing orthogonalization methods to estimate $\mathrm{msgn}(M)$. We implement Newton-Schulz iterations as provided in Muon (Jordan et al., ) with 5 iterations. For our low-rank orthogonalization methods, we set $r=0.1n$ as the input rank parameter, use command torch.linalg.qr to compute the Q factor, and apply 5 iterations of Newton-Schulz scheme following Muon (Jordan et al., ) to compute the matrix sign of $Q^{T}M$. In addition, we use command torch.pca_lowrank to efficient perform truncated SVD, setting the rank to $0.1n$ for truncation.
@@
We present a comparison of the computation time on GPUs for all competing methods in Figure 1, and the time distribution of our low-rank orthogonalization method with Gaussian sketching in Figure 2. From Figure 1, we observe that our low-rank orthogonalization method significantly reduces computation time compared to Newton-Schulz iterations and truncated SVD. Although both truncated SVD and our low-rank orthogonalization exploit low-rank structure for computation, truncated SVD is not well-suited for GPU environments and is therefore slower than the Newton-Schulz iterations. From Figure 2, we observe that in our low-rank orthogonalization method with Gaussian sketching, the QR decomposition accounts for approximately half of the total time, while the Newton-Schulz iterations and other operations (such as generating Gaussian random matrices and performing matrix multiplications) each make up roughly half of the remaining time.
Noise Robustness
@@
For each $n\in\{1000,2000,5000,10000\}$, we first randomly generate 10 nearly low-rank matrices $M\in\mathbb{R}^{n\times n}$ following strategy: the top $0.1n$ singular values are set to $1$, the remaining singular values are set to $10^{-4}$, and the singular vectors are randomly generated orthogonal vectors. For each $M$, we generate 50 noise matrices $N\in\mathbb{R}^{n\times n}$, with each entry drawn from a Gaussian distribution with mean zero and variance $\sigma^{2}\in\{0.1,1,10\}$, and construct noisy matrices $M_{N}=M+N$. We next apply Newton-Schulz iterations, and our low-rank orthogonalization method based on Gaussian sketching (Algorithm 1), using an input rank parameter $0.1n$, to estimate $\mathrm{msgn}(M_{N})$. For each $M$, we compute the trace of the empirical covariance matrix of the estimates of $\mathrm{msgn}(M_{N})$ produced by both Newton-Schulz iterations and Algorithm 1. Both methods are implemented in the same way as in the above synthetic experiments to illustrate the computational efficiency.
@@
Figure 3 presents a comparison of the average trace of the covariance matrices across all competing methods for matrix sign estimation. From this figure, we observe that, compared to applying the Newton-Schulz iterations to the full matrix, our low-rank orthogonalization method yields matrix sign estimates that are significantly less sensitive to noise. This is because singular vectors associated with small singular values are more sensitive to noise than others, so removing them improves the robustness of the matrix sign computation. As a result, our low-rank method provides more stable and robust estimates of the matrix sign function, particularly for nearly low-rank matrices.